Determine the interval of convergence for the power series

"\displaystyle\sum_{n=1}^ ∞" 3n(x-1)n/(n+1)2

1
Expert's answer
2021-11-17T03:47:03-0500

Let "a_n=\dfrac{3^n(x-1)^n}{(n+1)^2}." Then


"\bigg|\dfrac{a_{n+1}}{a_n}\bigg|=\bigg|\dfrac{\dfrac{3^{n+1}(x-1)^{n+1}}{(n+1+1)^2}}{\dfrac{3^n(x-1)^n}{(n+1)^2}}\bigg|"

"=3(\dfrac{n+1}{n+2})^2|x-1|\to3|x-1|\ as \ n\to\infin"

By the Ratio Test, the given series converges if "3|x-1|<1" and diverges if "3|x-1|>1."

Thus it converges if "|x-1|<\dfrac{1}{3}" and diverges if "|x-1|>\dfrac{1}{3}" .


"|x-1|<\dfrac{1}{3}"

"-\dfrac{1}{3}<x-1<\dfrac{1}{3}"

"\dfrac{2}{3}<x-1<\dfrac{4}{3}"

This means that the series converges in the interval "(\dfrac{2}{3}, \dfrac{4}{3})."

If "x=\dfrac{2}{3}," the series becomes


"\displaystyle\sum_{n=1}^{\infin}\dfrac{3^n(\dfrac{2}{3}-1)^n}{(n+1)^2}=\displaystyle\sum_{n=1}^{\infin}\dfrac{(-1)^n}{(n+1)^2}"

The series "\displaystyle\sum_{n=1}^{\infin}\dfrac{1}{(n+1)^2}=\displaystyle\sum_{k=2}^{\infin}\dfrac{1}{k^2}" converges as "p" -series with "p=2>1."

Then the series "\displaystyle\sum_{n=1}^{\infin}\dfrac{(-1)^n}{(n+1)^2}" converges absolutely.

If "x=\dfrac{4}{3}," the series becomes


"\displaystyle\sum_{n=1}^{\infin}\dfrac{3^n(\dfrac{4}{3}-1)^n}{(n+1)^2}=\displaystyle\sum_{n=1}^{\infin}\dfrac{1}{(n+1)^2}"

The series "\displaystyle\sum_{n=1}^{\infin}\dfrac{1}{(n+1)^2}=\displaystyle\sum_{k=2}^{\infin}\dfrac{1}{k^2}" converges as "p" -series with "p=2>1."

Therefore the interval of convergence is "\bigg[\dfrac{2}{3}, \dfrac{4}{3}\bigg]."